# Large-Scale Calculations of $\beta$-Decay Rates and Implications for $r$-Process Nucleosynthesis

### v1

@@ 3.1 Benchmarking theoretical rate calculations
As a first step in validating new rate calculations, it is important to compare the results with available data from NUBASE2020 (Kondev et al., 2021). Since the half-lives span orders of magnitude, it is common to employ the following quantifiers for the mean value:
@@
Results of the inter-model comparison are shown in Fig. 1, by restricting the summations in Eqs. (5) and (6) to nuclei with experimental half-lives below $T_{1/2}^{\rm exp,max}$, i.e., $T_{1/2}^{\text{exp.}(i)}<T_{1/2}^{\rm exp,max}$. Our DD-PC1 and DD-PCX results are compared to SkO’ (Ney et al., 2020), D3C* Marketin et al. (2016), and FRDM (Möller et al., 2003, 2019) predictions. In this comparison, we only consider nuclei with $T_{1/2}^{\rm calc.}<10^{6}$ s. As shown in Fig. 1(a), all models tend to overestimate the experimental half-lives. Both DD-PC1 and DD-PCX show a similar pattern, with the maximum discrepancy around $T_{1/2}^{\rm exp,max}=10$ s. In particular, DD-PCX shows excellent accuracy at $T_{1/2}^{\rm exp,max}=10^{4}$ s, with the mean relative error under 5%. For nuclei with shorter experimental half-lives, the maximum discrepancy is 20% for DD-PCX and 32% for DD-PC1. Overall, the mean error of our models is similar to other global theoretical models. In Fig. 1(b), we show the spread of the predictions around the mean, Eq. (6). Here, all models display a universal trend where $\sigma$ gradually increases with $T_{1/2}^{\rm exp,max}$.
@@
In Fig. 2 we compare different theoretical $\beta$-decay rates, across the nuclear landscape relative to DD-PC1. Overall, both EDFs from this work show excellent agreement with SkO’. Discrepancies are seen closer to the valley of stability, where half-lives are quite long, meaning that the rates are determined by only a few peaks of the strength function within the $Q_{\beta}$-window. Otherwise, regions of somewhat larger discrepancies exist around magic nuclei with $N=50,Z=28$, $N=82,Z=50$, and $N=184$. While our rates tend to be smaller than SkO’ for these nuclei, the differences are within an order of magnitude. The excellent agreement with SkO’ towards the neutron drip-line is particularly remarkable, considering significant model extrapolations involved. Additionally, both EDFs used in this work show very consistent predictions, indicating a relatively small systematic uncertainty due to the choice of a functional. On the other hand, the D3C* spherical model exhibits significant discrepancies for $Z>60$ and $N>126$. We note that it is difficult to assess whether the differences solely originate from the assumption of sphericity in D3C* or due to the specifics of effective interaction. Somewhat smaller differences are obtained for FRDM, around one order of magnitude, being especially pronounced around and above the $Z=82$ shell closure. The half-lives predicted in this work, $Q_{\beta}$ values, and other ground-state properties are available in a public repository (Saito et al., 2025).
@@
Figure 3: Final abundance patterns (top), the evolution of abundance-averaged proton number $\langle Z\rangle_{Y}$ as a function of time (middle), and the fission flow and neutron abundance as functions of time (bottom), for the s20ye0050 (left column) and s40ye0175 (right column) trajectories, respectively. The scaled solar abundance pattern (Goriely, 1999) is shown for reference.
3.2 Impact on the $r$-process
@@
We have performed nuclear reaction network calculations employing the new DD-PC1 and DD-PCX values of $\beta$-decay rates and the existing predictions of SkO’, FRDM, and D3C*. Overall, we find that the $r$-process calculations using the DD-PC1 and DD-PCX $\beta$-decay rates are generally consistent with the SkO’ calculations, and are significantly different from those with FRDM and D3C*.
@@
Since the results obtained with the nuclear properties taken from the REACLIB nuclear data library and from FRDM are very similar, in the following we will only show the results based on FRDM data in the main text. The complete set of results can be found in the Supplementary Material (Ravlić et al., 2025).
3.2.1 Parameterized trajectories
@@
We use ten parameterized trajectories to systematically assess the effect of different $\beta$-decay predictions. Our calculations show that $\beta$-decay rates significantly impact the abundances above the second $r$-process peak at $A\sim 130$. Therefore, to highlight this effect, we focus on trajectories corresponding to neutron-rich conditions s20ye0050 ($s/k_{B}=20$ and $Y_{e}=0.05$) and s40ye0175 ($s/k_{B}=40$ and $Y_{e}=0.175$), where $s/k_{B}$ and $Y_{e}$ are the initial entropy per baryon and electron fraction, respectively. s20ye0050 approximates an extremely neutron-rich condition, such as dynamical ejecta from a neutron star merger, where fission cycling is expected to occur, whereas s40ye0175 is a moderately neutron-rich condition. The complete set of results for ten trajectories can be found in Ravlić et al. (2025).
@@
In Fig. 3(a,b), most notable effects of $\beta$ decays on the final abundance patterns can be seen around $140\leq A\leq 170$ for s20ye0050 and $130\leq A\leq 180$ for s40ye0175, respectively. Fig. 3(c,d) shows that the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates consistently lead to a slower increase in the average proton number, compared to FRDM or D3C*. Since the neutron density of the environment decreases as the ejecta expand over time, the delayed population of fissile nuclei also delays the fission cycling process, primarily driven by neutron-induced fission and subsequent neutron capture and $\beta$ decay. In an extremely neutron-rich condition s20ye0050, a significant number of neutron-induced fission events still occur, as shown in Fig. 3(e). However, compared to D3C* and FRDM, subsequent neutron capture on the fission fragments would be hindered by the delayed timing of neutron-induced fission, leading to a dip in abundance patterns around $140\leq A\leq 170$ with DD-PC1, DD-PCX, and SkO’. In the case of s40ye0175, Fig. 3(f) shows that the delayed population of fissile nuclei leads to a significant decrease in the flow of neutron-induced fission. This results in a significant decrease in abundance below $A\sim 180$ due to the absence of fission cycling. While neutron-induced fission has been considered to be the dominant channel in the $r$-process nucleosynthesis (Panov et al., 2005; Martínez-Pinedo et al., 2007; Petermann et al., 2012), our calculations with the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates suggest that the flow of neutron-induced fission is comparable to, or only slightly larger than, that of $\beta$-delayed fission (see Ravlić et al. (2025)). Although for the fission yields we assume a 50/50 split, as the focus of the work is the new $\beta$-decay rates, in future studies, for a more detailed assessment of the effect of fission yields, we plan to obtain realistic fission yields using the consistent EDF approach (Flynn et al., 2022; Lay et al., 2024).
3.2.2 Trajectories from Astrophysical Simulations
@@
We consider two astrophysical simulations to provide more realistic $r$-process thermodynamic conditions: a mangeto-rotational supernova (Obergaulinger & Aloy, 2017; Reichert et al., 2021a), which we refer to as “MRSN”, and accretion disk outflow from a binary neutron star merger (Metzger & Fernández, 2014; Holmbeck et al., 2019), which we refer to as “disk outflow”. Each case consists of a set of ten thermodynamic trajectories, and the results in select trajectories will be highlighted. The complete results of the nuclear reaction network calculations are given in  Ravlić et al. (2025).
@@
MRSN has been studied as a site that can produce elements up to the third $r$-process peak in jet-like ejecta along the axis of the rotation under strong magnetic fields, by escaping the exposure to neutrinos, which would make the ejecta less neutron-rich (Mösta et al., 2014, 2015, 2018; Halevi & Mösta, 2018; Cameron, 2003; Reichert et al., 2022; Winteler et al., 2012; Nishimura et al., 2006, 2015, 2017; Symbalisty et al., 1985; Reichert et al., 2021a).
@@
Figure 4(a) shows the final abundance patterns for the most neutron-rich condition among the MRSN trajectories, and this is the only trajectory capable of synthesizing elements up to the third $r$-process peak. With the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates, the third $r$-process peak is reduced, compared to FRDM or D3C*. This indicates that $\beta$-decay rates, especially in the vicinity of and past the $N=126$ neutron shell closure, compete with the expansion timescale of the ejecta and substantially affect the nucleosynthetic outcome. The calculations with the REACLIB nuclear data library in Ravlić et al. (2025) yield more pronounced third $r$-process peaks, however, DD-PC1, DD-PCX, and SkO’ still consistently predict smaller peaks than FRDM and D3C*. These results suggest that conclusions on the production of elements in the third $r$-process peak and above in the MRSN conditions may be substantially dependent on nuclear physics inputs for the nucleosynthesis calculations, in particular $\beta$-decay rates.
@@
Holmbeck et al. (2019), who considered similar astrophysical conditions, have suggested that most actinide enrichments in metal-poor $r$-process enhanced stars are explained by an $r$-process source with a non-dominant amount of very neutron-rich, fission-cycling component. Given the significant effect of $\beta$ decay on fission flows indicated by our results, the understanding of actinide enrichment, in particular the distribution of the neutron-richness ($Y_{e}$) of the ejecta, is also likely affected by the choice of $\beta$-decay rates. Finally, the effect of $\beta$ decay on spontaneous fission flows is also present for the parameterized and disk outflow trajectories, consistent with the findings in Lund et al. (2023).
4 Conclusions

### v2

@@ III.1 Benchmarking theoretical rate calculations
As a first step in validating new rate calculations, it is important to compare the results with available data from NUBASE2020 [Nubase2020]. Since the half-lives span orders of magnitude, it is common to employ the following quantifiers for the mean value:
@@
Results of the inter-model comparison are shown in Fig. 1, by restricting the summations in Eqs. (5) and (6) to nuclei with experimental half-lives below $T_{1/2}^{\rm exp,max}$, i.e., $T_{1/2}^{\text{exp.}(i)}<T_{1/2}^{\rm exp,max}$. Our DD-PC1 and DD-PCX results are compared to SkO’ [Ney2020], D3C* [Marketin_RQRPA_2016], and FRDM [Moller2003, Moller2019] predictions. In this comparison, we only consider nuclei with $T_{1/2}^{\rm calc.}<10^{6}$ s. As shown in Fig. 1(a), all models tend to overestimate the experimental half-lives. Both DD-PC1 and DD-PCX show a similar pattern, with the maximum discrepancy around $T_{1/2}^{\rm exp,max}=10$ s. In particular, DD-PCX shows excellent accuracy at $T_{1/2}^{\rm exp,max}=10^{4}$ s, with the mean relative error under 5%. For nuclei with shorter experimental half-lives, the maximum discrepancy is 20% for DD-PCX and 32% for DD-PC1. Overall, the mean error of our models is similar to other global theoretical models. In Fig. 1(b), we show the spread of the predictions around the mean, Eq. (6). Here, all models display a universal trend where $\sigma$ gradually increases with $T_{1/2}^{\rm exp,max}$.
Figure 3: The residuals, $Q_{\beta}^{\rm calc}-Q_{\beta}^{\rm exp}$, between calculated and experimental $Q_{\beta}$-values for DD-PC1 (a) and DD-PCX (b) EDFs, as a function of experimental $Q_{\beta}^{\rm exp}$-values taken from AME2020 [Wang2021]. The dataset is split into even-even (+), odd-AA (circles), and odd-odd (x) nuclei. The solid line indicates the mean value, and the band shows the $1\sigma$ interval. Distribution of the residuals, along with the normal distribution fitted to the data, is shown in insets.
@@
In Fig. 2 we compare different theoretical $\beta$-decay rates, across the nuclear landscape relative to DD-PC1. Overall, both EDFs from this work show excellent agreement with SkO’. Discrepancies are seen closer to the valley of stability, where half-lives are quite long, meaning that the rates are determined by only a few peaks of the strength function within the $Q_{\beta}$ window. Otherwise, regions of somewhat larger discrepancies exist around magic nuclei with $N=50,Z=28$, $N=82,Z=50$, and $N=184$. While our rates tend to be smaller than SkO’ for these nuclei, the differences are within an order of magnitude. The excellent agreement with SkO’ towards the neutron drip line is particularly remarkable, considering significant model extrapolations involved. Additionally, both EDFs used in this work show very consistent predictions, indicating a relatively small systematic uncertainty due to the choice of a functional. On the other hand, the D3C* spherical model exhibits significant discrepancies for $Z>60$ and $N>126$. We note that it is difficult to assess whether the differences solely originate from the assumption of sphericity in D3C* or due to the specifics of effective interaction. Somewhat smaller differences are obtained for FRDM, around one order of magnitude, being especially pronounced around and above the $Z=82$ shell closure. The half-lives predicted in this work, $Q_{\beta}$ values, and other ground-state properties are available in a public repository [GitHub].
III.2 $Q_{\beta}$-value calculations
The $Q_{\beta}$-values are calculated by determining the relativistic Hartree-Bogoliubov ground states in the parent and daughter nuclei:
$Q_{\beta}^{\rm calc}=B(Z,N)-B(Z+1,N-1)+\Delta_{nH},$ (7)
where $B(Z,N)<0$ is the calculated binding energy and $\Delta_{nH}=0.782$ MeV is the neutron-hydrogen atom mass difference. Calculations of $Q_{\beta}$ values are challenging from a theoretical perspective because they require explicit considerations of odd-$A$ and odd-odd nuclei, which are often approximated by considering one- and two-quasiparticle excitations on top of even-even vacuum [Engel1999a]. In this work, we explicitly calculate binding energies of odd-$A$ and odd-odd nuclei within the EFA blocking procedure.
Figure 3 shows the residuals between the calculated and experimental $Q_{\beta}$ values, $Q_{\beta}^{\rm calc}-Q_{\beta}^{\rm exp}$, as a function of $Q_{\beta}^{\rm exp}$, for DD-PC1 (a) and DD-PCX (b) EDFs for 959 nuclei for which the experimental data are available [Wang2021]. The dataset is split into even-even, odd-$A$ and odd-odd parent nuclei. The mean value of the dataset and the $1\sigma$ deviation are marked. Regarding the global averages, the DD-PC1 calculations tend to slightly underestimate the experimental results with $\mu=-0.300$ MeV, whereas the DD-PCX calculations slightly overestimate the experiment with $\mu=0.157$ MeV. The $1\sigma$ discrepancy is also lower for DD-PCX by around 0.2 MeV, and it visibly produces a narrower distribution around the mean in Fig. 3(b). It is important to note that different classes of nuclei have different contributions to the total error. Since calculations of even-even or odd-odd nuclei require both proton and neutron blocking, either in the parent or in the daughter nucleus, it is to be expected that they are prone to larger errors. In particular, the mean value on the dataset of 233 even-even nuclei is $\mu=0.431(0.613)$ MeV for DD-PC1(DD-PCX) EDF, while for 260 odd-odd nuclei it is $\mu=-0.977(-0.280)$ MeV. On the other hand, odd-$A$ nuclei show the smallest mean deviations, being $\mu=-0.288(+0.173)$ MeV for DD-PC1(DD-PCX) EDF. The $1\sigma$ deviation for all the datasets does not exceed 1 MeV. Certainly, further improvements to EDFs, and especially concerning the isovector properties, are required to improve the accuracy of $Q_{\beta}$-values.
The insets in Fig. 3 show the distribution of the residuals. For both EDFs, the distributions are nearly-Gaussian, which is indicative of the absence of obvious systematic effects.
Figure 4: Correction procedure for half-lives to eliminate the spurious contribution of the negative strength for odd-$A$ and odd-odd nuclei. Panels (a) and (b) show the GT strength, decomposed into $K=0$ and $K=1$ components, together with their sum (Total) in 51Sc (a) and 52Sc (b). The vertical dashed line marks the manual correction made to remove the negative strength contribution, while the solid line denotes the limit (8). The result of the correction is shown in panel (c), resulting in a significant decrease of 51Sc and 52Sc half-lives. Calculations are performed by employing the DD-PC1 interaction.
III.3 Negative strength correction
As demonstrated in Ref. [Ney2020], contributions from $\beta^{+}$ strength can mix with $\beta^{-}$ strength once the EFA is employed within QRPA. The problem lies in the fact that for odd-$A$ and odd-odd nuclei, the ground-state energy of the daughter nucleus is unknown, and it has to be estimated to determine the lower end of the contour. In particular, based on the blocking calculation, the energy window, to the leading-order, can be estimated as:
$W_{0}^{\rm max}(m_{e}c^{2})\approx\Delta_{np}+\lambda_{n}-\lambda_{p},$ (8)
as obtained from Eq. (3). Although it works reasonably well for most nuclei and transitions, the above prescription can lead to the wrong choice of integration contour in odd-$A$ and odd-odd nuclei [cf. Eq. (1)]. Note that to obtain odd-even staggering one has to add $\Omega_{1}$ term to the above equation, $\Omega_{1}$ being the lowest QRPA eigenvalue. A prescription to estimate this quantity was offered in Ref. [Ney2020]. However, in this work, we neglect this term for the estimation of the contour, and manually correct the calculations for nuclei with unusually long half-lives. In Fig. 4 we show the GT strength function in 51Sc and 52Sc, explicitly indicating the contour limit using the prescription from Eq. (8), and then the corrected value (shift indicated by an arrow). In panel (c), we show the original half-lives calculated using the estimated window from Eq. (8) and the corrected values.
To systematically correct the half-lives, first, the “suspicious” nuclei are identified by inspecting the individual isotopic chains. We have identified approximately 30 nuclei for both DD-PC1 and DD-PCX interactions, which necessitated these corrections. We note that such calculations are numerically expensive, since they require explicit analysis of the strength function within the $Q_{\beta}$-window. To achieve the best possible resolution of individual peaks in the strength function, while keeping the computational time manageable, the smearing of the Lorentzian is set at $\eta=0.1$ MeV with energy spacing of $0.03$ MeV. Calculations are performed for the GT ($1^{+}$) as well as the FF ($1^{-}$) strength function.
III.4 Contribution of first-forbidden transitions
As described in Sec. II, in our calculations, we include both allowed GT ($1^{+}$) and FF ($0^{-},1^{-}$ and $2^{-}$) multipoles. Detailed treatment of FF transitions, including both matrix elements and contour integration of phase-space integrals is quite involving, and can be found in Ref. [Ravlic2025a]. It is interesting to study how the relative importance of FF transitions evolves across the nuclear landscape. In Fig. 5 we compare the contribution of FF transitions to the total $\beta$-decay rate using DD-PC1 and DD-PCX models, as well as the SkO’ model [Ney2020]. The universal trend in all models is that nuclei up to $Z\approx 70$ start, close to the valley of stability, with dominant contribution of GT strength to the total decay rate, which eventually decreases for more neutron-rich nuclei as the FF contribution increases. It is seen that up to $Z\approx 70$ and $N>82$, SkO’ calculations show a larger contribution of GT transitions for neutron-rich nuclei, whereas both DD-PC1 and DD-PCX estimate it to be around 40–60%. For $Z>70$, FF transitions tend to dominate, which is especially pronounced around the $N=126$ and $N=184$ shell closures, and an overall better agreement between the three models is obtained. (See also Refs. [Borzov2003, Robin2024] for the competition between allowed and first-forbidden transitions in $r$-process nuclei.) Both models from this work, DD-PC1 and DD-PCX tend to give very consistent predictions, except for the heaviest parts of the nuclide chart beyond $Z>90$ and $N>200$, where DD-PC1 predicts somewhat stronger FF contribution.
Figure 5: Contribution of first-forbidden (FF) transition (in %) for DD-PC1 (a) and DD-PCX (b) predictions of this work, and SkO’ results (c) from Ref. [Ney2020].
@@ III.4 Contribution of first-forbidden transitions
Figure 6: Final abundance patterns (a,b), the time-evolution of abundance-averaged proton number $\langle Z\rangle_{Y}$ (c,d), the neutron-induced fission flow and neutron abundance (e,f), and the $\beta$-delayed fission flow compared to the neutron-induced flow (g,h), for the s20ye0050 (left column) and s40ye0175 (right column) trajectories, respectively. The scaled solar abundance pattern from Ref. [Goriely1999_solar] is shown for reference.
III.5 Impact on the $r$-process
@@
We have performed nuclear reaction network calculations employing the new DD-PC1 and DD-PCX values of $\beta$-decay rates and the existing predictions of SkO’, FRDM, and D3C* models. Overall, we find that the $r$-process calculations using the DD-PC1 and DD-PCX $\beta$-decay rates are generally consistent with the SkO’ calculations, and are significantly different from those with FRDM and D3C*.
@@
Since the results obtained with the nuclear properties taken from the REACLIB nuclear data library and from FRDM are very similar, in the following we will only show the results based on FRDM data in the main text. The complete set of results can be found in Ref. [GitHub].
III.5.1 Parameterized trajectories
@@
We used ten parameterized trajectories to systematically assess the effect of different $\beta$-decay predictions. Our calculations show that $\beta$-decay rates significantly impact the abundances above the second $r$-process peak at $A\sim 130$. Therefore, to highlight this effect, we focus on trajectories corresponding to neutron-rich conditions s20ye0050 ($s/k_{B}=20$ and $Y_{e}=0.05$) and s40ye0175 ($s/k_{B}=40$ and $Y_{e}=0.175$), where $s/k_{B}$ and $Y_{e}$ are the initial entropy per baryon and electron fraction, respectively. s20ye0050 approximates an extremely neutron-rich condition, such as dynamical ejecta from a neutron star merger, where fission cycling is expected to occur, whereas s40ye0175 is a moderately neutron-rich condition. The complete set of results for ten trajectories can be found in a public repository [GitHub].
@@
In Fig. 6(a,b), most notable effects of $\beta$ decays on the final abundance patterns can be seen around $140\leq A\leq 170$ for s20ye0050 and $130\leq A\leq 180$ for s40ye0175, respectively. Fig. 6(c,d) shows that the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates consistently lead to a slower increase in the average proton number, compared to FRDM or D3C*. Since the neutron density of the environment decreases as the ejecta expand over time, the delayed population of fissile nuclei also delays the fission cycling process, primarily driven by neutron-induced fission and subsequent neutron capture and $\beta$ decay. In an extremely neutron-rich condition s20ye0050, a significant number of neutron-induced fission events still occur, as shown in Fig. 6(e). However, compared to D3C* and FRDM, subsequent neutron capture on the fission fragments would be hindered by the delayed timing of neutron-induced fission, leading to a dip in abundance patterns around $140\leq A\leq 170$ with DD-PC1, DD-PCX, and SkO’. In the case of s40ye0175, Fig. 6(f) shows that the delayed population of fissile nuclei leads to a significant decrease in the flow of neutron-induced fission. This results in a significant decrease in abundance below $A\sim 180$ due to the absence of fission cycling. While neutron-induced fission has been considered to be the dominant channel in the $r$-process nucleosynthesis [Panov2005, Martinezpinedo2007, Petermann2012], our calculations with the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates suggest that the flow of neutron-induced fission is comparable to, or only slightly larger than, that of $\beta$-delayed fission, as displayed in Fig. 6(g,h). Although for the fission yields we assume a 50/50 split, as the focus of the work is the new $\beta$-decay rates, in future studies, for a more detailed assessment of the effect of fission yields, we plan to obtain realistic fission yields using the consistent EDF approach developed in Refs. [Flynn2022, Lay2024].
III.5.2 Trajectories from Astrophysical Simulations
@@
We considered two astrophysical simulations to provide more realistic $r$-process thermodynamic conditions: a mangeto-rotational supernova [Obergaulinger2017, Reichert2021], which we refer to as “MRSN”, and accretion disk outflow from a binary neutron star merger [Metzger2014, Holmbeck2019], which we refer to as “disk outflow”. Each case consists of a set of ten thermodynamic trajectories, and the results in select trajectories will be highlighted. The complete results of the nuclear reaction network calculations are given in Ref. [GitHub].
@@
MRSN has been studied as a site that can produce elements up to the third $r$-process peak in jet-like ejecta along the axis of the rotation under strong magnetic fields, by escaping the exposure to neutrinos, which would make the ejecta less neutron-rich [Mosta_MRSN_2014, Mosta_MRSN_2015, Mosta_MRSN_2018, Halevi_MRSN_2018, Cameron_MRSN_2003, Reichert_MRSN_2022, Winteler_MRSN_2012, Nishimura_MRSN_2006, Nishimura_MRSN_2015, Nishimura_MRSN_2017, Symbalisty_MRSN_1985, Reichert2021].
@@
Figure 7(a) shows the final abundance patterns for the most neutron-rich condition among the MRSN trajectories, and this is the only trajectory capable of synthesizing elements up to the third $r$-process peak. With the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates, the third $r$-process peak is reduced, compared to FRDM or D3C*. This indicates that $\beta$-decay rates, especially in the vicinity of and past the $N=126$ neutron shell closure, compete with the expansion timescale of the ejecta and substantially affect the nucleosynthetic outcome. The calculations with the REACLIB nuclear data library yield more pronounced third $r$-process peaks, however, DD-PC1, DD-PCX, and SkO’ still consistently predict smaller peaks than FRDM and D3C* [GitHub]. These results suggest that conclusions on the production of elements in the third $r$-process peak and above in the MRSN conditions may be substantially dependent on nuclear physics inputs for the nucleosynthesis calculations, in particular $\beta$-decay rates.
@@
In Ref. [Holmbeck2019], considering similar astrophysical conditions, it was suggested that most actinide enrichments in metal-poor $r$-process enhanced stars are explained by an $r$-process source with a non-dominant amount of very neutron-rich, fission-cycling component. Given the significant effect of $\beta$ decay on fission flows indicated by our results, the understanding of actinide enrichment, in particular the distribution of the neutron-richness ($Y_{e}$) of the ejecta, is also likely affected by the choice of $\beta$-decay rates. Finally, the effect of $\beta$ decay on spontaneous fission flows is also present for the parameterized and disk outflow trajectories, consistent with the findings in Ref. [Lund2023].
IV Conclusions